Stickybuns:

In which direction must the graph of f(x) = x be shifted to produce the graph of g(x) = x + 7? A. up and right B. down C. up D. left and down

Stickybuns:

either a or c?

Stickybuns:

@Angle

Nnesha:

https://www.desmos.com/calculator type in both equation y=x and y=x+7

Nnesha:

y=mx+b where b represent the vertical shift if b is positive =up If b is negative = down

Stickybuns:

so down?

Nnesha:

g(x) = x + 7 what is b in this equation ?

Stickybuns:

positive

Nnesha:

that's right.

Stickybuns:

so positive = up

Nnesha:

yes.

Stickybuns:

c?

Nnesha:

yes